OpenStudy (anonymous):

Hey guys can you give multiple ways to solve this ?! integrate With respext to x : 1/(1+(x^4))

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ 1+ x^4 } dx\]

OpenStudy (anonymous):

Residues would be the best and simplest way for this..

OpenStudy (anonymous):

... if it were a definite integral, sure. Try partial fractions: \[x^4+1=-(x^2-\sqrt2x+1)(x^2+\sqrt2x+1)\] Then \[\begin{align*}-\frac{1}{(x^2-\sqrt2x+1)(x^2+\sqrt2x+1)}&=\frac{Ax+B}{x^2-\sqrt2x+1}+\frac{Cx+D}{x^2+\sqrt2x+1}\\\\ -1&=Ax^3+Bx^2+\sqrt2Ax^2+\sqrt2Bx+Ax+B\\ &~~~~~~~~~~~+Cx^3+Dx^2-\sqrt2Cx^2-\sqrt2Dx+Cx+D\\\\ -1&=(A+C)x^3+(\sqrt2A+B-\sqrt2C+D)x^2\\ &~~~~~~~~~~~+(A+\sqrt2B+C-\sqrt2D)x+B+D \end{align*}\] So you have \[\begin{cases}\begin{align*}A+C&=0\\\sqrt2A+B-\sqrt2C+D&=0\\A+\sqrt2B+C-\sqrt2D&=0\\ B+D&=-1\end{align*}\end{cases}\] Solutions for the constants (5th column of matrix, top to bottom gives A,B,C,D): http://www.wolframalpha.com/input/?i=rref+%7B%7B1%2C0%2C1%2C0%2C0%7D%2C%7BSqrt%5B2%5D%2C1%2C-Sqrt%5B2%5D%2C1%2C0%7D%2C%7B1%2CSqrt%5B2%5D%2C1%2C-Sqrt%5B2%5D%2C0%7D%2C%7B0%2C1%2C0%2C1%2C-1%7D%7D

OpenStudy (anonymous):

So now you're integrating \[\frac{1}{2\sqrt2}\int\frac{x-\sqrt2}{x^2-\sqrt2x+1}~dx-\frac{1}{2\sqrt2}\int\frac{x+\sqrt2}{x^2+\sqrt2x+1}~dx\]

OpenStudy (anonymous):

thanks guys for your helpful repiles ! If someone gets another way ... just type it !! You guys stay awesome!!